OpenStudy (anonymous):

find an equation of the tangent line to the graph of f at the given point. f(x)=x√x^2+5) point (2,6)

OpenStudy (anonymous):

i have to use the product rule, right?

OpenStudy (anonymous):

Nah you got to use me

OpenStudy (anonymous):

a trinity rule? o.0

OpenStudy (anonymous):

y = x √(x^2 + 5) => dy/dx = √(x^2 + 5) + x * x / √(x^2 + 5) => dy/dx at (2, 6) = √(2^2 + 5) + 2^2 / √(2^2 + 5) = 3 + 4/3 = 13/3 The equation of the tangent line passes through (2, 6) and has a slope = 13/3 => the eqn. of the tangent line is y - 6 = (13/3) (x - 2) => 13x - 3y = 8.

OpenStudy (anonymous):

Trinity does rule =)

OpenStudy (anonymous):

that was fast.

OpenStudy (anonymous):

Like I said,I am trinity

OpenStudy (anonymous):

Need other help?

OpenStudy (anonymous):

in a few minutes, yea. just reading over your answer

OpenStudy (anonymous):

No doubt

OpenStudy (anonymous):

so how did you know to get the equation, => dy/dx = √(x^2 + 5) + x * x / √(x^2 + 5)

OpenStudy (anonymous):

You find the square root f:(X²+ 5))

ganeshie8 (ganeshie8):

product rule i think

OpenStudy (anonymous):

Post your other question?

OpenStudy (anonymous):

I might fall asleep waiting lol

OpenStudy (anonymous):

oh sorry, lol. these questions you can;t help me with :( i have to look at these graphs. thanks though. goodnight!

OpenStudy (anonymous):

;o